# A Kaczmarz-inspired approach to accelerate the optimization of neural network wavefunctions

### v1

@@ 4 Results
We now turn to our numerical results. Our main finding is that SPRING outperforms KFAC, MinSR, and MinSR+M for several small atoms and molecules. We present results for the carbon, nitrogen, and oxygen atoms in Section 4.1 and for the $N_{2}$ molecule in Section 4.2. We use VMCNet vmcnet2021github to run all our experiments, which is a neural network VMC framework based on JAX jax2018github . For all of our experiments, we use a standard FermiNet architecture with 16 dense determinants; see Table 1 for details.
@@
For our training phase we use one hundred thousand optimization iterations with one thousand MCMC samples per iteration. For our MCMC proposals we use Gaussian all-electrons moves with a step-size that is tuned dynamically to maintain an acceptance ratio of approximately 50%. We take ten MCMC steps between optimization iterations to reduce correlations between the samples, and we use mean–centered local energy clipping to improve the stability of the optimization. We use a decaying learning rate with a decay rate of $r=1e-4$ and learning rate schedule $\eta_{k}=\frac{\eta}{1+rk}$. For MinSR, MinSR+M, and SPRING the learning rate $\eta$ is equivalent to the imaginary time step $\delta\tau$. For more details on our VMC settings, see Table 2. To obtain final energies we run a separate inference phase with the parameters from the last iteration of each optimization run; see Table 3 for details.
@@
To compare the methods fairly, we do not start them from fully random initializations. Rather, for each system in question, we first run a short preliminary optimization and save the result to a checkpoint. The rest of our experiments then load the starting parameters from the end of the preliminary optimization phase, switch out the optimizer and hyperparameters to the desired settings, and run one hundred thousand further optimization iterations. This procedure removes the chaotic early optimization stage as a factor from the comparison between the different methods, ensuring that no method is unfairly advantaged by randomly having a favorable start. Our preliminary optimization is different from the Hartree-Fock pretraining procedure (see pfau2020ab ; von2022self for example) because we optimize the variational energy from the beginning. However, our results should translate well to a setting with pretraining as a result of this preliminary optimization. See Table 4 for more details on the preliminary optimization phase.
@@
For all of our optimizers we use a Tikhonov regularization parameter of $\lambda=0.001$ and a norm constraint of $C=0.001$ unless indicated otherwise. For MinSR+M, we use a momentum parameter of $\mu=0.9$, and for SPRING we use the regularization parameter $\mu=0.99$. In Section 4.3 we present several hyperparameter studies which demonstrate that these values provide a fair basis for comparing the methods.
@@
We note that we are using a somewhat smaller network, fewer MCMC samples, and fewer optimization iterations than were used in most state-of-the-art calculations with KFAC; hence it should not be surprising that some of our results do not reach the same level of accuracy as those presented in the literature for comparable systems. Our results can be viewed as as proof-of-principle for the value of SPRING on small calculations, which will need to be extended to larger calculations by future works.
@@ 4.1 Small atoms
We first present results for KFAC, MinSR, MinSR+M, and SPRING on the carbon, nitrogen, and oxygen atoms. We find that it is critical to turn the learning rate of each method in order to achieve the best performance and obtain a fair comparison. Thus, for each method, we test a number of learning rates on the carbon atom, with results presented in Figure 1. For the energy and variance trajectories, in this and all other figures, we smooth out the curves by reporting averages over a sliding window of ten thousand iterations. Our experiments lead us to choose a learning rate of $\eta=0.02$ for KFAC, $\eta=0.1$ for MinSR, $\eta=0.2$ for MinSR+M, and $\eta=0.02$ for SPRING.
@@
4.2 $N_{2}$ molecule
We next test KFAC, MinSR, MinSR+M, and SPRING on a somewhat larger and more challenging system: the $N_{2}$ molecule. For this system, we tune the learning rates at the equilibrium bond distance of 2.016 Bohr. We present the results of the learning rate sweeps in Figure 3. Based on these results, we pick learning rates of $\eta=0.05$ for KFAC, $\eta=0.02$ for MinSR, $\eta=0.02$ for MinSR+M, and $\eta=0.002$ for SPRING. These learning rates are not always chosen strictly for the best final energy since the final energy can vary due to statistical fluctuations in the parameters and statistical errors in the energy estimation. We instead look at the optimization trajectories and the final energies and pick the learning rate that appears to be optimal based on the combination of the two. It is worth noting that the optimal learning rates for MinSR+M and SPRING change by a factor of $10$ between the carbon atom and the $N_{2}$ molecule, whereas the optimal learning rate for MinSR changes by only a factor of $5$, and the optimal learning rate for KFAC changes by only a factor of $2$. The optimal learning rate may thus be more system-dependent for MinSR+M and SPRING compared to the other methods.
@@ 4.2 $N_{2}$ molecule
Using these optimized learning rates, we then compare the four methods both at equilibrium and at a stretched bond distance of 4.0 Bohr. The results of these two comparisons are presented in Figure 4. For these systems none of the optimizers are able to reach chemical accuracy, which is expected due to the relatively small network, few MCMC samples, and few optimization iterations that we are using. Nonetheless, SPRING still displays a marked advantage over the other methods, converging much faster and to a lower energy and energy variance in both cases. The MinSR+M optimizer only improves upon MinSR marginally for this system, and it is outperformed by KFAC. This provides evidence that the SPRING algorithm is distinctly better than simply adding momentum on top of MinSR.
@@ 4.3 Hyperparameter studies
In this section, we provide several experiments that demonstrate that our main results are robust to the choices of $C$ and $\mu$. For simplicity we focus on the carbon atom for the purposes of this section.
@@
Next, we test the MinSR+M scheme with several different values of the momentum parameter $\mu$, using the previously optimized learning rate of $\eta=0.2$. For values larger than $\mu=0.95$ we find that the optimization is unstable and we do not report numerical results. We find that choosing $\mu$ anywhere between $0.8$ and $0.95$ produces approximately optimal results, which justifies our use of $\mu=0.9$ for our main experiments. Results are shown in Figure 7(a). Finally, we test SPRING with several values of its regularization parameter $\mu$, again using the previously optimized learning rate of $\eta=0.02$. Results are shown in Figure 7(b). We find that choosing $\mu$ anywhere between $0.98$ and $0.999$ produces approximately optimal results, which justifies our choice of $\mu=0.99$ for our main experiments. Interestingly, if $\mu=1.0$ then the method becomes unstable. We do not yet have an explanation for the source of this instability.
@@ 4.4 Computational Cost
The computational cost of SPRING is essentially identical to that of MinSR, as the extra computations for SPRING are negligible in cost. The same applies to MinSR+M. For all three of these methods, the computational bottleneck lies in computing the matrix $T=\bar{O}\bar{O}^{T}$, which has asymptotic complexity $O(N_{p}\cdot N_{s}^{2})$. This cost dominates the $O(N_{s}^{3})$ cost of calculating the Cholesky factorization of $T$.
@@ 5 Discussion
Neural network wavefunctions represent a promising avenue towards highly accurate simulations of small but challenging molecular systems. The major bottleneck of applying such wavefunctions is the high cost associated with their optimization. In this work we introduce a new optimizer called SPRING to alleviate this bottleneck. SPRING combines ideas from the recently proposed MinSR optimizer minsr with ideas from the randomized Kaczmarz method for solving overdetermined least-squares problems strohmer2009randomized . By doing so, SPRING is able to utilize optimization history to improve upon MinSR in a principled way at essentially no extra cost. We test SPRING on several small atoms and molecules, comparing it against MinSR, MinSR with momentum, and KFAC. We find that SPRING consistently outperforms the alternatives across all tested systems. We hope that these findings will be extended to larger systems by future works.
@@
Due to several discrepancies between the VMC setting and the traditional Kaczmarz setting, we are not yet able to furnish a rigorous proof of convergence for the SPRING optimizer. One interesting direction for future research is to modify SPRING in a way that makes it possible to prove its convergence. For example, the exponentially convergent randomized Kaczmarz method of Strohmer and Vershynin requires that the rows are sampled with probability proportional to the square of their Euclidean norm strohmer2009randomized . Furthermore, in the randomized block Kaczmarz method, Needell and Tropp find that for optimal performance it is critical to partition the constraints of the least-squares problem into well-conditioned blocks needell2014paved . Such sampling schemes would represent a significant departure from traditional VMC methods that sample from the probability density of the wavefunction, but they could lead to rigorous convergence guarantees, better performance, or both. There are also variants of the Kaczmarz method such as the randomized extended Kaczmarz algorithm of Zouzias and Freris zouzias2013randomized that can converge to the solution of inconsistent least-squares problems. This technique cannot be directly applied to the VMC setting since it requires sampling the columns as well as the rows of the system. However, it could serve as inspiration for further development of the algorithm.
@@
We also draw a connection between the MinSR and SPRING methods for optimizing neural network wavefunctions and the efficient subsampled natural gradient descent method of Ren and Goldfarb ren2019efficient . In particular, we show in A that MinSR can be viewed as a simplified implementation of the method of Ren and Goldfarb which applies as long as the gradient of the loss function is a linear combination of the model gradients at the sampled points. This is quite a common scenario in machine learning, occurring for example in supervised learning with a mean-squared error loss function. In such a setting, SPRING can then be viewed as a potential improvement to existing subsampled natural gradient descent methods. We leave it to future works to determine whether SPRING can yield performance improvements for applications outside of VMC.
6 Acknowledgements
@@
This research used the Savio computational cluster resource provided by the Berkeley Research Computing program at the University of California, Berkeley (supported by the UC Berkeley Chancellor, Vice Chancellor for Research, and Chief Information Officer). This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number(s) DE-SC0023112 (G.G.). This effort was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Number N00014-23-1-2729 (N.A.). LL is a Simons Investigator in Mathematics. We thank Yixiao Chen, Zhiyan Ding, Yuehaw Khoo, Michael Lindsey, Eric Neuscamman, and Zaiwen Wen for their helpful discussions.
7 Disclaimer

### v2

@@ 4 Results
We now turn to our numerical results. Our main finding is that SPRING outperforms KFAC, MinSR, and MinSR+M for several small atoms and molecules. We present results for the carbon, nitrogen, and oxygen atoms in Section 4.1 and for the $N_{2}$ and CO molecules in Section 4.2. We summarize the final energies attained in our main experiments in Table 1. We use VMCNet [30] to run all our experiments, which is a neural network VMC framework based on JAX [39]. For all of our experiments, we use a standard FermiNet architecture with 16 dense determinants; see Table 2 for details.
@@
For our training phase we use one hundred thousand optimization iterations with one thousand MCMC samples per iteration. For our MCMC proposals we use Gaussian all-electrons moves with a step-size that is tuned dynamically to maintain an acceptance ratio of approximately 50%. We take ten MCMC steps between optimization iterations to reduce correlations between the samples, and we use mean–centered local energy clipping to improve the stability of the optimization. We use a decaying learning rate with a decay rate of $r=1e-4$ and learning rate schedule $\eta_{k}=\frac{\eta}{1+rk}$. For MinSR, MINSR+M, and SPRING, we assume $\delta\tau=1$ and apply the learning rate $\eta_{k}$ as described in Section 3. For more details on our VMC settings, see Table 3. To obtain final energies we run a separate inference phase with the parameters from the last iteration of each optimization run; see Table 4 for details.
@@
To compare the methods fairly, we do not start them from fully random initializations. Rather, for each system in question, we first run a short preliminary optimization and save the result to a checkpoint. The rest of our experiments then load the starting parameters from the end of the preliminary optimization phase, switch out the optimizer and hyperparameters to the desired settings, and run one hundred thousand further optimization iterations. This procedure removes the chaotic early optimization stage as a factor from the comparison between the different methods, ensuring that no method is unfairly advantaged by randomly having a favorable start. Our preliminary optimization is different from the Hartree-Fock pretraining procedure (see [3, 6] for example) because we optimize the variational energy from the beginning. However, our results should translate well to a setting with pretraining as a result of this preliminary optimization. See Table 5 for more details on the preliminary optimization phase.
@@
For all of our optimizers we use a Tikhonov regularization parameter of $\lambda=0.001$ and a norm constraint of $C=0.001$ unless indicated otherwise. For MinSR+M we use a momentum parameter of $\mu=0.9$, and for SPRING we use the regularization parameter $\mu=0.99$. In Section 4.3 we present several hyperparameter studies which demonstrate that these values provide a fair basis for comparing the methods. We also investigate the effect of the preliminary optimization phase and demonstrate that SPRING outperforms KFAC significantly without preliminary optimization.
@@
In all of our numerical results we report energy errors relative to the benchmark energies listed in Table 1. For the energy and variance trajectories, we smooth out the curves by reporting averages over a sliding window of ten thousand iterations. We note that we are using a somewhat smaller network, fewer MCMC samples, and fewer optimization iterations than were used in most state-of-the-art calculations with KFAC; hence it should not be surprising that some of our results do not reach the same level of accuracy as those presented in the literature for comparable systems. Our results can be viewed as as proof-of-principle for the value of SPRING on small calculations, which will need to be extended to larger calculations by future works.
System KFAC MinSR MinSR+M SPRING Benchmark Benchmark source
C -37.8445 -37.8445 -37.8448 -37.8449 -37.8450 [40], Table XI
N -54.5877 -54.5885 -54.5889 -54.5890 -54.5892 [40], Table XI
O -75.0643 -75.0652 -75.0659 -75.0668 -75.0673 [40], Table XI
N2, equilibrium -109.5301 -109.5268 -108.5294 -109.5322 -109.5423 [41], Table II
[…]
CO -113.3140 -113.3092 -113.3117 -113.3169 -113.3255 [3], Table II
Table 1: Summary of final energies attained with tuned learning rates for all systems studied. Energies are reported in Hartrees, to four decimal places.
@@ 4.1 Small atoms
We first present results for KFAC, MinSR, MinSR+M, and SPRING on the carbon, nitrogen, and oxygen atoms. We find that it is critical to turn the learning rate of each method in order to achieve the best performance and obtain a fair comparison. Thus, for each method, we test a number of learning rates on the carbon atom, with results presented in Figure 1. Our experiments lead us to choose a learning rate of $\eta=0.02$ for KFAC, $\eta=0.1$ for MinSR, $\eta=0.2$ for MinSR+M, and $\eta=0.02$ for SPRING.
@@
4.2 Molecules
We next test KFAC, MinSR, MinSR+M, and SPRING on two somewhat larger systems: the N2 and CO molecules. We begin by tuning the learning rates for all four methods on the N2 molecule at an equilibrium bond distance of 2.016 Bohr. We present the results of the learning rate sweeps in Figure 3. Based on these results, we pick learning rates of $\eta=0.05$ for KFAC, $\eta=0.02$ for MinSR, $\eta=0.02$ for MinSR+M, and $\eta=0.002$ for SPRING. These learning rates are not always chosen strictly for the best final energy since the final energy can vary due to statistical fluctuations in the parameters and statistical errors in the energy estimation. We instead look at the optimization trajectories and the final energies and pick the learning rate that appears to be optimal based on the combination of the two. It is worth noting that the optimal learning rates for MinSR+M and SPRING change by a factor of $10$ between the carbon atom and the N2 molecule, whereas the optimal learning rate for MinSR changes by only a factor of $5$, and the optimal learning rate for KFAC changes by only a factor of $2$. The optimal learning rate may thus be more system-dependent for MinSR+M and SPRING compared to the other methods.
@@ 4.2 Molecules
Using these optimized learning rates, we then compare the four methods on the N2 molecule both at equilibrium and at a stretched bond distance of 4.0 Bohr, and on the CO molecule at an equilibrium bond distance of 2.173 Bohr. The results of these comparisons are presented in Figure 4 and Figure 5. For these systems none of the optimizers are able to reach chemical accuracy, which is expected due to the relatively small network, few MCMC samples, and few optimization iterations that we are using. Nonetheless, SPRING still displays a marked advantage over the other methods, converging much faster and to a lower energy and energy variance in all three cases. The MinSR+M optimizer only improves upon MinSR marginally for these systems, and it is outperformed by KFAC. This provides evidence that the SPRING algorithm is distinctly better than simply adding momentum on top of MinSR.
@@
Figure 5: Comparison of methods on CO molecule, with learning rates tuned on N2.
@@ 4.3 Hyperparameter studies
In this section, we provide several experiments that demonstrate that our main results are robust to the choices of $C$ and $\mu$. We also provide evidence that the results are not significantly skewed by the preliminary optimization phase. For simplicity we focus on the carbon atom for the purposes of this section.
@@
To better understand the impact of the norm constraint on SPRING, we return to the learning rate sweep for SPRING on the carbon atom and provide some additional data from these five optimization runs. In particular, we plot the scale factor $q_{k}=\max(1,\left\lVert\eta_{k}\phi_{k}\right\rVert/\sqrt{C})$. For example, if $q_{k}=5$ then this means that at iteration $k$ the norm constraint resulted in a scaling down of the parameter update by a factor of $5$, and if $q_{k}=1$ then the norm constraint at iteration $k$ had no effect. We see in Figure 9 that $q_{k}$ tends to grow over the course of the first one hundred optimization iterations and then decay to $1$ as the optimization progresses. We can understand this in the following way: $q_{k}$ first rises as more optimization history is incorporated into SPRING; it then decays due to both the learning rate decay and the decay of the energy gradient. As should be expected, with a larger learning rate, $q_{k}$ reaches a larger peak and takes longer to decay to $1$.
@@
Figure 9: Effect of the norm constraint on the SPRING parameter update during the learning rate sweep on the carbon atom. The quantity $q_{k}=\max(1,\left\lVert\eta_{k}\phi_{k}\right\rVert/\sqrt{C})$ represents the extent to which the parameter update is scaled down as a result of the norm constraint.
Next, we test the MinSR+M scheme with several different values of the momentum parameter $\mu$, using the previously optimized learning rate of $\eta=0.2$. For values larger than $\mu=0.95$ we find that the optimization is unstable and we do not report numerical results. We find that choosing $\mu$ anywhere between $0.8$ and $0.95$ produces approximately optimal results, which justifies our use of $\mu=0.9$ for our main experiments. Results are shown in Figure 10(a). Similarly, we test SPRING with several values of its regularization parameter $\mu$, again using the previously optimized learning rate of $\eta=0.02$. Results are shown in Figure 10(b). We find that choosing $\mu$ anywhere between $0.98$ and $0.999$ produces approximately optimal results, which justifies our choice of $\mu=0.99$ for our main experiments. Interestingly, if $\mu=1.0$ then the method becomes unstable. We do not yet have an explanation for the source of this instability.
Finally, we consider the effects of the preliminary optimization phase. Since we have used KFAC for the preliminary optimization in all cases, it is unclear thus far whether SPRING can perform well when run from a random initialization. To ameliorate this concern, we compare the performance of SPRING against KFAC without any preliminary optimization, with results in Figure 11. We find that KFAC outperforms SPRING in the very early stages of the optimization. However, SPRING overtakes KFAC after about seven thousand training iterations and reaches chemical accuracy in fewer than half as many iterations as KFAC. We conclude that SPRING is also effective starting from a random initialization.
@@
Figure 11: Comparison of SPRING versus KFAC on the carbon atom without preliminary optimization. We use the previously tuned learning rate of $\eta=0.02$ for both KFAC and SPRING.
@@ 4.4 Computational Cost
The computational cost of SPRING is essentially identical to that of MinSR, as the extra computations for SPRING are negligible in cost. The same applies to MinSR+M. For all three of these methods, the computational bottleneck lies in computing the matrix $T=\bar{O}\bar{O}^{T}$, which has asymptotic complexity $O(N_{p}\cdot N_{s}^{2})$. This cost dominates the $O(N_{s}^{3})$ cost of the Cholesky factorization.
@@ 5 Discussion
Neural network wavefunctions represent a promising avenue towards highly accurate simulations of small but challenging molecular systems. The major bottleneck of applying such wavefunctions is the high cost associated with their optimization. In this work we introduce a new optimizer called SPRING to alleviate this bottleneck. SPRING combines ideas from the recently proposed MinSR optimizer [26] with ideas from the randomized Kaczmarz method for solving overdetermined least-squares problems [28]. By doing so, SPRING is able to utilize optimization history to improve upon MinSR in a principled way at essentially no extra cost. We test SPRING on several small atoms and molecules, comparing it against MinSR, MinSR with momentum, and KFAC. We find that SPRING consistently outperforms the alternatives across all tested systems. We hope that these findings will be extended to larger systems by future works.
@@
Due to several discrepancies between the VMC setting and the traditional Kaczmarz setting, we are not yet able to furnish a rigorous proof of convergence for the SPRING optimizer. One interesting direction for future research is to modify SPRING in a way that makes it possible to prove its convergence. For example, the exponentially convergent randomized Kaczmarz method of Strohmer and Vershynin requires that the rows are sampled with probability proportional to the square of their Euclidean norm [28]. Furthermore, in the randomized block Kaczmarz method, Needell and Tropp find that for optimal performance it is critical to partition the constraints of the least-squares problem into well-conditioned blocks [29]. Such sampling schemes would represent a significant departure from traditional VMC methods that sample from the probability density of the wavefunction, but they could lead to rigorous convergence guarantees, better performance, or both. There are also variants of the Kaczmarz method such as the randomized extended Kaczmarz algorithm of Zouzias and Freris [43] that can converge to the solution of inconsistent least-squares problems. This technique cannot be directly applied to the VMC setting since it requires sampling the columns as well as the rows of the system. However, it could serve as inspiration for further development of the algorithm.
@@
We also draw a connection between the MinSR and SPRING methods for optimizing neural network wavefunctions and the efficient subsampled natural gradient descent method of Ren and Goldfarb [25]. In particular, we show in A that MinSR can be viewed as a simplified implementation of the method of Ren and Goldfarb which applies as long as the gradient of the loss function is a linear combination of the model gradients at the sampled points. This is quite a common scenario in machine learning, occurring for example in supervised learning with a mean-squared error loss function. In such a setting, SPRING can then be viewed as a potential improvement to existing subsampled natural gradient descent methods. We leave it to future works to determine whether SPRING can yield performance improvements for applications outside of VMC.
Acknowledgements
@@
This research used the Savio computational cluster resource provided by the Berkeley Research Computing program at the University of California, Berkeley (supported by the UC Berkeley Chancellor, Vice Chancellor for Research, and Chief Information Officer). This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number(s) DE-SC0023112 (G.G.). This effort was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Number N00014-23-1-2729 (N.A.). LL is a Simons Investigator in Mathematics. We thank Yixiao Chen, Zhiyan Ding, Yuehaw Khoo, Michael Lindsey, Eric Neuscamman, and Zaiwen Wen for their helpful discussions, and the anonymous reviewers for their valuable comments.
Disclaimer